A ball is thrown 125 meters upward and then falls the same distance back on earth. Neglecting air resistance, its total time in the air is


1
Expert's answer
2021-09-21T15:03:39-0400

Solution.

"h=125m;"

"h=\dfrac{gt^2}{2};"

"t^2=\dfrac{2h}{g};"

"t=\sqrt{\dfrac{2h}{g}};"

"t=\sqrt{\dfrac{2\sdot125}{10}}=5s;"

The fall time is equal to the rise time to the same height.

Answer: 10s.